OpenStudy (alfordabby):

A motorboat travels 150mi in 6 hours going upstream. It travels 282mi going downstream in the same amount of time. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (alfordabby):

150 / 6 = 25

OpenStudy (alfordabby):

282 / 6 = 47

OpenStudy (alfordabby):

2s = 72 s = 36

OpenStudy (alfordabby):

36 - c = 25 c = 11